OpenStudy (anonymous):

Can Someone Help Plsss! Traveling at 55 miles per hour, how many minutes, rounded to the nearest whole number, does it take to drive 310 miles from Boston to Syracuse?

OpenStudy (texaschic101):

1 hr = 60 minutes now we will set up a proportion 55/60 = 310/x --- 55 miles to 60 minutes = 310 miles to x minutes cross multiply (55)(x) = (60)(310) 55x = 18600 x = 18600/55 x = 338 2/11 to the nearest whole number is 338